Q1: The momentum of a body increases by 20%. What is the % increase in it`s K.E .Q2: two vectors of magnitude 3 & 4 have resultant 5, the angle between vectors is ?Please solve these questions for me. I have Medical University Test on 21st Oct.

Grade:12

1 Answers

Arun
6 years ago

Kinetic Energy=(1/2)m(v^2) ;Momentum(p)=mv ;

Kinetic Energy={ (mv)^2}/2m={(p)^2}/2m

where m=Mass of a body and v=Velocity

Momentum=p

Increase in Momentum=20%

So New Momentum (p')=p * {1+ (20/100)} =1.2 p

New Kinetic energy={(p')^2}/2m

(p')^2= (1.2 p)^2=1.44 p

As mass of a body is constant

Therefore % Increase in K.E.= 44%(1.44p-p=0.44p)

 

Now cor 2nd question,

As 3,4 and 5 makes a Pythagorean triplet,

That means angle between 3 and 4 unit vectors is 90 degree